# PositionIC: Unified Position and Identity Consistency for Image Customization

### v4

@@ Abstract
Recent subject-driven image customization excels in fidelity, yet fine-grained instance-level spatial control remains an elusive challenge, hindering real-world applications. This limitation stems from two factors: a scarcity of scalable, position-annotated datasets, and the entanglement of identity and layout by global attention mechanisms. To this end, we introduce PositionIC, a unified framework for high-fidelity, spatially controllable multi-subject customization. First, we present BMPDS, the first automatic data-synthesis pipeline for position-annotated multi-subject datasets, effectively providing crucial spatial supervision. Second, we design a lightweight, layout-aware diffusion framework that integrates a novel visibility-aware attention mechanism. This mechanism explicitly models spatial relationships via an NeRF-inspired volumetric weight regulation to effectively decouple instance-level spatial embeddings from semantic identity features, enabling precise, occlusion-aware placement of multiple subjects. Extensive experiments demonstrate PositionIC achieves state-of-the-art performance on public benchmarks, setting new records for spatial precision and identity consistency. Our work represents a significant step towards truly controllable, high-fidelity image customization in multi-entity scenarios. Code and data will be publicly released.
@@ 1 Introduction
Diffusion-based models have recently revolutionized visual synthesis, especially in text-to-image (T2I) generation, where they produce photorealistic images closely aligned with short textual prompts [sdxl, stablediffusion3, ldm, playground, flux]. Beyond generic generation, subject-driven image customization aims to integrate one or more reference objects into user-specified scenes under the guidance of a text prompt. Traditionally, this task emphasizes two core goals: (i) preserving the appearance identity of the reference object(s), and (ii) aligning the synthesis with the user’s textual description, where remarkable improvements in visual fidelity have been achieved by recent methods [uno, dreamo, omini_control, wang2025unicombine].
@@
Despite substantial progress in identity-preserving personalization methods, current approaches still struggle with flexible placement and layout control. Specifically, two coupled bottlenecks remain: (1) the scarcity of large-scale datasets annotated with explicit positional labels of multiple subjects, hindering the learning of spatial reasoning; and (2) the prevalent use of global-level attention mechanisms in diffusion transformers, which inadvertently entangle semantic identity and spatial layout, thereby failing to provide fine-grained, visibility-aware control. While a few recent works [ms-diffusion, gligen] have attempted layout control, they typically exhibit a trade-off between spatial precision and subject identity consistency, thus limiting their practical utility.
@@
Specifically, to overcome the data bottleneck, we devise an automatic pipeline that expands single-subject collections into scalable, high-quality multi-subject datasets annotated with explicit position masks. A hierarchical training schedule establishes a bidirectional generation paradigm, progressively moving from single-to-multi subject synthesis and back, so that resolution constraints are relaxed while subject drift is suppressed. Because of the inherent hallucination of Multi-modal Large Language Models (MLLMs) [han2024instinctive, bpo], we avoid direct visual comparisons: expert vision models first translate visual content into textual descriptions, after which MLLMs perform multi-dimensional consistency checks. This two-stage filtering markedly improves data reliability.
Built upon the BMPDS corpus, PositionIC directly integrates our visibility-aware attention mechanism into the diffusion transformer architecture. This mechanism is achieved mainly via an operation inspired by NeRF [nerf] termed Volumetric Weight Regulation, which explicitly constructs attention masks based on the likelihood of visibility derived from volumetric rendering principles. Unlike prevalent global-level attention, our visibility-aware attention mechanism effectively decouples instance-level spatial embeddings from semantic identity features. This design enables independent and accurate placement of each subject by explicitly modeling occlusion and perspective relationships, and crucially introduces no extra train-time parameters or inference overhead. Restricting reference features to user-specified regions further enhances identity fidelity and spatial precision, unlocking new applications of controllable image customization (Figure 1).
@@ 2.1 Subject-driven Generation
In addition to using text prompts for conditional image generation, current diffusion models [ELITE, li2023blip, huang2024realcustom, ip-adapter, groundingbooth, fastcomposer, MultiDiffusion, feng2025personalize] support reference image input to achieve preservation of subject identity. Dreambooth [dreambooth] and LoRa [lora] control the generation of diffusion models through fine-tuning on the specific subject. Recently, Diffisuion-Transformers-based subject-driven models [uno, omini_control, omini_control2, dreamo, xiao2025omnigen, flux-kontext] further advance subject-driven generation. They introduce reference images as context information through token concatenation to ensure the consistency of objects during generation.
@@ 2.2 Position Controllable Generation
Some works [instancediffusion, regional_flux, wang2024compositional, jimenez2023mixture, MultiDiffusion, shi2025layoutcot, LaRender] have attempted to generate with precise layout control. Gligen [gligen] encodes Fourier embedding as grounding tokens to inject position information. MS-diffusion [ms-diffusion] utilizes a grounding resampler correlating visual information with specific entities and spatial constraints. However, they still suffer from inconsistencies in vision and position, which hinder further application.
@@ 3.1 Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS)
Position-controllable image-driven customization requires high-fidelity paired data, featuring prominent subjects and high resolution with layout control signals. However, existing open-source datasets such as Subject200K [omini_control] generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information. We introduce BMPDS, a Bidirectional Multi-dimensional Perception Data Synthesis framework to tackle these limitations. We adopt a hierarchical generation-and-selection strategy to progressively improve data quality, gradually introducing single-image and multi-image pairs with spatial control information during generation. The overall framework is depicted in Figure 2.
@@ 3.1.2 Multi-dimensional Perception Data Filter
Previous studies [cao2024mmfuser, he2025analyzing] have shown that MLLMs have limited capability in recognizing fine-grained details in images. Instead of directly feeding image pairs into the MLLMs for filtering, we establish a reliable system to achieve more accurate and efficient data filtering. Specifically, we divide the filtering process into three levels based on granularity. Firstly, we segment the subjects from data pairs and utilize CLIP-I [CLIP] and DINO [DINO] scores $s_{v}$ to filter out images with significantly lower consistency. Subsequently, we pass two subjects’ images through MLLMs (e.g., GPT-4o), which directly gives a similarity score $s_{vlm}$ based on shape, color and so on. Lastly, we employ Describe Anything Model (DAM) [describe_anything], a description expert to obtain detailed textual descriptions for each subject. Given these textual description pairs, we instruct GPT-4o as a judge to autonomously select comparative features(e.g., color, shape) and assign multi-dimensional similarity scores $s_{ds}$.
@@ 3.2 PositionIC
When extending text-to-image models to image-to-image tasks, the fidelity of the generated object largely depends on the magnitude of the attention between the corresponding regions in the target image and the reference image. We refer to this as attention accumulation. As shown in Figure 3, the attention map can be divided into four areas: text-text self attention, image-image self attention, text-image and image-text cross attention. If extending to single or multi-subject custom generation via concatenating method like [uno, omini_control, flux-kontext, eligen], the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
To achieve effective attention accumulation, we propose Visibility-Aware Attention, which utilizes volume rendering(VR) to explicitly define the area that can be focused on for each reference image. Previous works [eligen, regional_flux, zhang2025creatidesign, LaRender] explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT). The overall mechanism is shown in Figure 3.
@@ 3.2.1 Volume Rendering(VR)
The weighted mask utilizes VR [volume_render] to handle the occlusion relationships between individual objects. Similar to NeRF [nerf], the expected color $C(r)$ observed by the camera along the distance from the far bound $t_{f}$ to the near bound $t_{n}$ is:
@@ 4.1.2 Evaluation Metrics
For subject-generation tasks, we evaluate subject similarity using CLIP-I, DINO-I on Dreambench [dreambooth]. For text fidelity, we calculate CLIP-T scores, which measure cosine similarity between the text embedding and image embedding from CLIP.
For the position-guided task, we evaluate different methods on our proposed PositionIC-Bench. We use Vision-R1 [visionr1] to determine the bounding box of the subject and calculate mIoU and AP scores with the label.
@@
RPF [regional_flux] 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
MS-Diffusion [ms-diffusion] 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
Instance-Diffusion [instancediffusion] 0.789 0.593 / 0.683 / 0.632 0.799 0.497 / 0.699 / 0.546
Gligen [gligen] 0.808 0.632 / 0.865 / 0.811 0.825 0.628 / 0.858 / 0.811

### v6

@@ Abstract
Recent subject-driven image customization excels in fidelity, yet fine-grained instance-level spatial control remains an elusive challenge, hindering real-world applications. This limitation stems from two factors: a scarcity of scalable, position-annotated datasets, and the entanglement of identity and layout by global attention mechanisms. To this end, we introduce PositionIC, a unified framework for high-fidelity, spatially controllable multi-subject customization. First, we present BMPDS, the first automatic data-synthesis pipeline for position-annotated multi-subject datasets, effectively providing crucial spatial supervision. Second, we design a lightweight, layout-aware diffusion framework that integrates a novel visibility-aware attention mechanism. This mechanism explicitly models spatial relationships via an NeRF-inspired volumetric weight regulation to effectively decouple instance-level spatial embeddings from semantic identity features, enabling precise, occlusion-aware placement of multiple subjects. Extensive experiments demonstrate PositionIC achieves state-of-the-art performance on public benchmarks, setting new records for spatial precision and identity consistency. Our work represents a significant step towards truly controllable, high-fidelity image customization in multi-entity scenarios. Code and data: https://github.com/MeiGen-AI/PositionIC.
@@ 1 Introduction
Diffusion-based models have recently revolutionized visual synthesis, especially in text-to-image (T2I) generation, where they produce photorealistic images closely aligned with short textual prompts [23, 5, 25, 17, 14]. Beyond generic generation, subject-driven image customization aims to integrate one or more reference objects into user-specified scenes under the guidance of a text prompt. Traditionally, this task emphasizes two core goals: (i) preserving the appearance identity of the reference object(s), and (ii) aligning the synthesis with the user’s textual description, where remarkable improvements in visual fidelity have been achieved by recent methods [35, 21, 28, 30].
@@
Despite substantial progress in identity-preserving personalization methods, current approaches still struggle with flexible placement and layout control. Specifically, two coupled bottlenecks remain: (1) the scarcity of large-scale datasets annotated with explicit positional labels of multiple subjects, hindering the learning of spatial reasoning; and (2) the prevalent use of global-level attention mechanisms in diffusion transformers, which inadvertently entangle semantic identity and spatial layout, thereby failing to provide fine-grained, visibility-aware control. While a few recent works [33, 18] have attempted layout control, they typically exhibit a trade-off between spatial precision and subject identity consistency, thus limiting their practical utility.
@@
Specifically, to overcome the data bottleneck, we devise an automatic pipeline that expands single-subject collections into scalable, high-quality multi-subject datasets annotated with explicit position masks. A hierarchical training schedule establishes a bidirectional generation paradigm, progressively moving from single-to-multi subject synthesis and back, so that resolution constraints are relaxed while subject drift is suppressed. Because of the inherent hallucination of Multi-modal Large Language Models (MLLMs) [7, 22], we avoid direct visual comparisons: expert vision models first translate visual content into textual descriptions, after which MLLMs perform multi-dimensional consistency checks. This two-stage filtering markedly improves data reliability.
Built upon the BMPDS corpus, PositionIC directly integrates our visibility-aware attention mechanism into the diffusion transformer architecture. This mechanism is achieved mainly via an operation inspired by NeRF [20] termed Volumetric Weight Regulation, which explicitly constructs attention masks based on the likelihood of visibility derived from volumetric rendering principles. Unlike prevalent global-level attention, our visibility-aware attention mechanism effectively decouples instance-level spatial embeddings from semantic identity features. This design enables independent and accurate placement of each subject by explicitly modeling occlusion and perspective relationships, and crucially introduces no extra train-time parameters or inference overhead. Restricting reference features to user-specified regions further enhances identity fidelity and spatial precision, unlocking new applications of controllable image customization (Figure 1).
@@ 2.1 Subject-driven Generation
In addition to using text prompts for conditional image generation, current diffusion models [34, 16, 10, 39, 38, 36, 1, 6] support reference image input to achieve preservation of subject identity. Dreambooth [26] and LoRa [9] control the generation of diffusion models through fine-tuning on the specific subject. Recently, Diffisuion-Transformers-based subject-driven models [35, 28, 29, 21, 37, 15] further advance subject-driven generation. They introduce reference images as context information through token concatenation to ensure the consistency of objects during generation.
@@ 2.2 Position Controllable Generation
Some works [32, 4, 31, 12, 1, 27, 40] have attempted to generate with precise layout control. Gligen [18] encodes Fourier embedding as grounding tokens to inject position information. MS-diffusion [33] utilizes a grounding resampler correlating visual information with specific entities and spatial constraints. However, they still suffer from inconsistencies in vision and position, which hinder further application.
@@ 3.1 Bidirectional Multi-dimensional Perception Data Synthesis (BMPDS)
Position-controllable image-driven customization requires high-fidelity paired data, featuring prominent subjects and high resolution with layout control signals. However, existing open-source datasets such as Subject200K [28] generate paired data using diptych images, which leads to object inconsistency issues and are limited by low resolution and the lack of positional information. We introduce BMPDS, a Bidirectional Multi-dimensional Perception Data Synthesis framework to tackle these limitations. We adopt a hierarchical generation-and-selection strategy to progressively improve data quality, gradually introducing single-image and multi-image pairs with spatial control information during generation. The overall framework is depicted in Figure 2.
@@ 3.1.2 Multi-dimensional Perception Data Filter
Previous studies [2, 8] have shown that MLLMs have limited capability in recognizing fine-grained details in images. Instead of directly feeding image pairs into the MLLMs for filtering, we establish a reliable system to achieve more accurate and efficient data filtering. Specifically, we divide the filtering process into three levels based on granularity. Firstly, we segment the subjects from data pairs and utilize CLIP-I [24] and DINO [3] scores $s_{v}$ to filter out images with significantly lower consistency. Subsequently, we pass two subjects’ images through MLLMs (e.g., GPT-4o), which directly gives a similarity score $s_{vlm}$ based on shape, color and so on. Lastly, we employ Describe Anything Model (DAM) [19], a description expert to obtain detailed textual descriptions for each subject. Given these textual description pairs, we instruct GPT-4o as a judge to autonomously select comparative features(e.g., color, shape) and assign multi-dimensional similarity scores $s_{ds}$.
@@ 3.2 PositionIC
When extending text-to-image models to image-to-image tasks, the fidelity of the generated object largely depends on the magnitude of the attention between the corresponding regions in the target image and the reference image. We refer to this as attention accumulation. As shown in Figure 3, the attention map can be divided into four areas: text-text self attention, image-image self attention, text-image and image-text cross attention. If extending to single or multi-subject custom generation via concatenating method like [35, 28, 15, 41], the attention map will expand to nearly three times its original size, making it harder for the model to focus on the corresponding regions.
To achieve effective attention accumulation, we propose Visibility-Aware Attention, which utilizes volume rendering(VR) to explicitly define the area that can be focused on for each reference image. Previous works [41, 4, 42, 40] explore the effort of restricting the attention horizon between special words and noise, we extend it to subject-driven generation task to unlock the positional control ability of diffusion transformers (DiT). The overall mechanism is shown in Figure 3.
@@ 3.2.1 Volume Rendering(VR)
The weighted mask utilizes VR [13] to handle the occlusion relationships between individual objects. Similar to NeRF [20], the expected color $C(r)$ observed by the camera along the distance from the far bound $t_{f}$ to the near bound $t_{n}$ is:
@@ 4.1.2 Evaluation Metrics
For subject-generation tasks, we evaluate subject similarity using CLIP-I, DINO-I on Dreambench [26]. For text fidelity, we calculate CLIP-T scores, which measure cosine similarity between the text embedding and image embedding from CLIP.
For the position-guided task, we evaluate different methods on our proposed PositionIC-Bench. We use Vision-R1 [11] to determine the bounding box of the subject and calculate mIoU and AP scores with the label.
@@
RPF [4] 0.341 0.015 / 0.063 / 0.007 0.369 0.070 / 0.002 / 0.011
MS-Diffusion [33] 0.501 0.097 / 0.329 / 0.075 0.421 0.028 / 0.146 / 0.005
Instance-Diffusion [32] 0.789 0.593 / 0.683 / 0.632 0.799 0.497 / 0.699 / 0.546
Gligen [18] 0.808 0.632 / 0.865 / 0.811 0.825 0.628 / 0.858 / 0.811